

********************************************************************
UARS WEEKLY STATUS REPORT - 28 APR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct for new
telemetry requirements, PEM will need to update processing code with a patch
delivery. Updated HEPS calibration parameters were finalized and are ready for
delivery. Updated MEPS quality calibration files were finalized and are ready
for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell, have
been fixed; his suggestions have been incorporated with in the PEM production
code. Both updated HEPS and MEPS calibration data, along with the Level 3
code changes, will influence the in-situ energy deposition predictions, so new
versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to operate normally.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues with no
problems. The MPG continues to assist the Project Scientist with the scheduling
of instrument operations.
MPG provided support to the ACRIM instrument representative in transferring the
May Daily Activity Plans to the CMS database.
***********
CDHF
***********
STATUS
__________
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day SEN (5.0) PROD (1992) (1R-1C=0L)
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, MLS,
SUSIM, Solstice, and WINDII.
- Continue to test HRDI V4.4 on the Rehost System.
- Preparing to test ISAMS V12.0 and PEM V3.32 software on the
Rehost System.
- Verifying correlative ingests NMC, ORBIT etc.
- Verifying LEVEL0, Quicklook and NRT. Successfully processed
Attitude extraction, Observatory Log, and Sensor Extraction Software
against this data on the Rehost System and found no differences on what
was processed on the VAX Cluster.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 April, 2000 (Orbit 46968, GMT Day 106) through 21 April
2000 (Orbit 47072, GMT Day 112).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 111/0226 47044 Off
The beta angle ranged between the angles of 40.8 degrees to 17.7 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+1.0 degrees on 25 April 2000 (DOY 116).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 14 April 2000 and 21 April 2000 are provided in Appendix D.
The clock error ranged from -11.0 to 5.7 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 02
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 19
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/111 show the observatory data loss to be 295 hours, 06
minutes, 29.145 seconds* (an increase of 0 minutes 0 seconds since the last
report period).
*These calculations have been recalculated to include 1999/346 through 2000/006
statistics using scheduled times.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
There were more problems encountered with the UTTS this week. Both the FOT and
its support staff are doing everything possible to correct these problems so
that the testing of the new safehold TMONs can begin.
The FOT continues to develop an online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure has
yet to be finalized and must undergo peer review before it will be ready.
MEETINGS
The FOT attended the UARS SERB meeting on 19 April 2000. The next SERB meeting
is scheduled for 26 April 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for April 25, 2000 (DOY 116). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 April 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46972 106/0715 00-163 NBTR NBTR OPS DOY 106
46979 106/1820 00-165 NBTR NBTR OPS DOY 107
46984 107/0218 00-166 OBC Clk Adj 20.6 to 0.00 msec/day
46987 107/0801 00-165 NBTR NBTR OPS DOY 107
46995 107/1958 00-167 NBTR NBTR OPS DOY 108
47000 108/0358 00-167 NBTR NBTR OPS DOY 108
47003 108/0805 00-167 NBTR NBTR OPS DOY 108
47009 108/1823 00-168 NBTR NBTR OPS DOY 109
47017 109/0725 00-168 NBTR NBTR OPS DOY 109
47024 109/1745 00-169 NBTR NBTR OPS DOY 110
47031 110/0543 00-169 NBTR NBTR OPS DOY 110
47035 110/1243 00-169 NBTR NBTR OPS DOY 110
47038 110/1748 00-170 NBTR NBTR OPS DOY 111
47044 111/0226 00-171 WINDII POWER OFF, LOW S/C PWR
47048 111/0810 00-170 NBTR NBTR OPS DOY 111
47049 111/1111 00-170 NBTR NBTR OPS DOY 111
47055 111/2008 00-171 NBTR NBTR OPS DOY 112
47060 112/0406 00-172 NBTR NBTR OPS DOY 112
47061 112/0549 00-172 NBTR NBTR OPS DOY 112
47069 112/1754 00-173 NBTR NBTR OPS DOY 113
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 April 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 46985-7,47060-3
WIQBVOLTAGE Y-LO WINDII Quiet Bus Voltage 46986
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 April 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46984 107/02:21 (15 Apr 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 April 2000
14 April 2000 - GMT Day 105 Beta = 40.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ---------- ----------- ------ ----- ------
1 -33.6/-50.4 -3.05/-4.31 +0.0/+0.0 15.8 0.000 0.0
2 +89.6/+0.0 +5.79/+4.51 +28.0/-15.2 24.5 1.040 20.9
3 +22.4/-44.8 +5.15/+3.55 +27.6/-15.2 24.5 1.038 20.6
21 April 2000 - GMT Day 112 Beta = 17.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ---------- ---------- ------- ------- ------
1 -22.4/-44.8 -3.05/-4.31 +0.0/+0.0 15.8 0.000 0.0
2 +84.0/-5.6 +6.44/+5.15 +30.4/-14.4 24.0 1.030 22.5
3 +33.6/-44.8 +5.15/+3.87 +29.6/-15.2 23.8 1.030 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 April 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-014 FOT 46992 107/1559 22235 PTP CRASHED
00-015 FOT 47016 109/0540 N/A TAC 6 CRASHED
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-011 FOT 46932 103/1550 N/A BLG 32 POWER OUTAGE
00-013 FOT 46946 104/1410 N/A TAC 4 CRASH
00-014 FOT 46992 107/1559 22235 PTP CRASHED
00-015 FOT 47016 109/0540 N/A TAC 6 CRASHED
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list April 00 14170 Generic Late Acq.
SEVNTFW00/16